OpenStudy (anonymous):

Solve: 3(3t + 3) = 3(-3t + -39)

OpenStudy (anonymous):

distribute out the 3's in both sides then it should be simpler or cancel out the 3 by dividing both sides by 3. THis one is easier I think.

OpenStudy (anonymous):

9t+9=-9t - 117. 18t + 9 = -117. 18t=-126. t=-126/18 t=-7

OpenStudy (anonymous):

or if you cancaled out the 3s 3t+3=-3t+-39